Suppose a population is known to be normally distributed with a mean, μ, equal to 144 and a standard deviation, σ, equal to 27. Approximately what percent of the population would be between 144 and 171?
Answer #1

Solution :

Given that ,

mean = = 144

standard deviation = =27

P(144< x <171 ) = P[(144-144) / 27< (x - ) / < (171 -144) /27 )]

= P( 0< Z < 1)

= P(Z < 1) - P(Z < 0)

Using z table   

= 0.8413-0.1587

=0.6826

=68.26%